two dice, one red and one green are tossed, find
p(red die even score/green die even score)
1
Expert's answer
2012-10-08T09:51:40-0400
There are

N = 3² = 9

results when both red and green diсe have even score (they are 2:2,2:4,2:6, 4:2,4:4,4:6, 6:2,6:4 and 6:6). The total number of possible events is

T = 6² = 36.

So, the probability

p(red die even score/green die even score) = N/T = 9/36 = 0.25.